determine the sum for the positive two-digit integers that are not divisible by 5
Answer by fractalier(2101)

The easiest way to do this is to sum up all the two digit numbers and subtract those divisible by 5...
The sum of the numbers 10 to 99 = (90/2)(10+99) = 4905
The sum of the numbers 10, 15, 20, ..., 95 = (20/2)(10+95) = 1050
4905 - 1050 =
3855